# Holstein polarons, Rashba-like spin splitting and Ising superconductivity in electron-doped MoSe2

### v1

@@ I Introduction
The resulting polaronic bound state can be described by Fröhlich or Holstein Hamiltonians. The Fröhlich model describes the interaction of electrons with optical phonons, which leads to the formation of large polarons, whose radii are much larger than the lattice constant. The Holstein model, on the other hand, describes the interaction with acoustic phonons, which leads to the formation of small polarons, whose radii are of the same order of magnitude as the lattice constant Emin (2013).
@@
Figure 1(a) shows the multi-valley scattering between $\Sigma$ and ${\rm K}$ valleys in the Brillouin zone. The electron scattering vectors $q_{1}$ and $q_{2}$ correspond to either $\Gamma$-${\rm K}$ or $\Gamma$-${\rm M}$ phonon wave vectors. For both MoS2 and MoSe2 the acoustic phonon modes have similar energies for ${\rm K}$ and ${\rm M}$ and the dominant contribution to electron-phonon coupling arises from the highest energy acoustic phonon mode. This behaviour has been demonstrated both experimentally Zeng et al. (2012) and theoretically Kaasbjerg et al. (2012) for MoS2. However, only ${\rm K}$ valley Holstein polarons (Fig. 1(b)) have been observed in the ARPES study of Rb-dosed MoS2 surface  Kang et al. (2018). Second, the spin-orbit coupling (SOC) should induce the splitting of the electron bands at $\Sigma$ and $\Sigma$’ valleys due to inversion symmetry breaking on the surface with alkali metal dosing. However, in all previous ARPES studies of alkali-dosed semiconducting TMDCs the spin-split features at $\Sigma$ valley have not been resolved Zhang et al. (2014); Eknapakul et al. (2014); Alidoust et al. (2014); Riley et al. (2015); Miwa et al. (2015); Zhang et al. (2016); Kim et al. (2016, 2017); Kang et al. (2017, 2018); Han et al. (2021).
Figure 2: Evolution of the MoSe2 electronic structure with Rb dosing: (a-a5) ARPES spectra of $\Gamma$-${\rm K}$ conduction band dispersions measured near the Fermi energy for dosing 0, 20, 60, 100, 220, 580s. (b-b5) Corresponding valence band spectra in a wide energy range.
The isostructural MoSe2 compound is the perfect model system for probing the multi-valley physics induced by electron-doping. Density functional theory (DFT) calculations show that in contrast to MoS2, the conduction band minimum (CBM) in bulk MoSe2 is at the $\Sigma$, rather than at ${\rm K}$ valley (see Supplementary Fig. S1). For the alkali-metal dosing on MoSe2 surface one expects to have larger SOC band-splitting as in the case of MoS2. The accessibility of the $\Sigma$ valley, as well as the ${\rm K}$ valley expected at higher doping Kim et al. (2016, 2017), makes MoSe2 the best candidate to study physics of multi-valley polaron modes using high-resolution ARPES.
@@ II.1 Evolution of the electronic band structure with dosing
In our ARPES data shown in Fig. 2 the spectral intensity appears at the Fermi level, marking the occupation of the conduction band and corresponding semiconductor to conductor transition with the Rb dosing on the MoSe2 surface. After initial dosing the maximum spectral weight at the Fermi level is observed around $\Sigma$ valley, showing an indirect band gap $E_{g}^{\Gamma-\Sigma}$ $\sim$1.4 eV. However, with increasing amount of dosed Rb the relative intensity of $\Sigma$ valley decreases, whereas the intensity of ${\rm K}$ valley increases. Correspondingly, it appears that the conduction band minimum shifts from $\Sigma$ to ${\rm K}$ valley as in case of monolayer MoSe2 (see Fig. 2(a1-a3)). Withing observed energy broadening the band minima at $\Sigma$ and ${\rm K}$ are very close (see Supplementary Fig. S5), incidentally making electron doped MoSe2 perfect system for studying multi-valley physics. Simultaneously with the shift of the CBM intensity, the intensity of the valence band compared to the conduction band decreases (see Fig. 2 (b1-b3)). This first crossover at around 40 s of Rb dosing marks the transition from the electron doped bulk to the more “monolayer-like” doped surface conduction band structure Eknapakul et al. (2014); Riley et al. (2015); Kang et al. (2017); Kim et al. (2017).
@@ II.2 Holstein polarons in “surface doping” regime
The measured Fermi surface shows two electron-like pockets at $\Sigma$ valley and single electron-like pocket at ${\rm K}$ valley. The second derivative images of $\Sigma$ and ${\rm K}$ valleys in Fig. 3(c,d) both show well-defined discontinuities of the spectral function separated by 20 meV. Moreover, the photoemission spectra integrated the momentum range between $\Sigma$ and ${\rm K}$ valleys in Fig. 5(a), show discrete non-dispersing energy levels with the same energy separation (dashed vertical black lines in Fig. 5(b)).
@@ II.3 Spin-splitting at the $\Sigma$ valley in “surface doping” regime
While polarons at the spin-degenerate ${\rm K}$-valley has been observed previously for MoS2 Kang et al. (2018), polarons at the spin-polarised $\Sigma$ valleys are discovered for the first time. The spin-orbit splitting of the electron band dispersion at $\Sigma$ valley is apparent in ARPES data measured along the $\Gamma$-${\rm K}$ direction. Figure 4 shows band dispersions at the $\Sigma$ valley together with the Rashba-like model fit constrained by the band positions from the corresponding momentum distribution curves (MDCs) at the Fermi level.
@@
The increase of the Fermi surface size indicates the gradual electron doping of MoSe2 with Rb dosing. From the same data one could see that the magnitude of band-splitting is also increasing with electron doping.This Rashba-like spin-splitting of the bands comes from inversion symmetry breaking due to the electric field created by electron charge transfer from dosed alkali atoms at the surface. Such inversion symmetry breaking on the surface has been seen at the valence band of semiconducting Kang et al. (2017); Kim et al. (2017) as well as at the conduction band of semimetallic Clark et al. (2019) alkali dosed TMDC materials. Figure 4(f) shows how the magnitude of the observed splitting consistently and monotonically becomes larger as the surface charge increases. This observation aligns perfectly with our understanding that the band splitting is Rashba-like and arises from the combination of spin-orbit coupling with the increasing inversion symmetry-breaking electric field at the surface.
@@ III Conclusion
By studying electronic structure evolution of MoSe2 with Rb surface dosing we have identified three different regimes: surface doping, inter-layer intercalation and ordered alkali metal layer formation. Figure 6(a) shows the two dimensional (2D) electron carrier density as a function of the alkali metal dosing time, calculated using the Luttinger’s theorem and experimentally determined sizes of the Fermi surface. In the “surface doping” regime we observe continuous electron doping of MoSe2 with corresponding increase of the Fermi surface size and value of the Rashba-like spin-orbit splitting.
@@
While our results show that the increasing surface electric field description is no longer applicable when we dose Rb atoms above 120 s, we could extrapolate a gradual increase of the 2D electron carrier density with the alkali metal dosing time within the ”surface doping” regime up to the superconducting transition. In this case, using our ARPES data we discover that maximum Tc doping level corresponds to the Fermi surface topology where the electron-like pocket at ${\rm K}$ valley touches the outer electron pocket at $\Sigma$ valley (see Fig. 6(b)). In another words, the maximum Tc corresponds to the Fermi level being located exactly at the conduction band saddle point between ${\rm K}$ and $\Sigma$ valleys. This can lead to a dramatic enhancement of pairing strength, due to strong increase of the momentum resolved electron-phonon coupling constant $\lambda$, like in case of Rb intercalated PtTe2  Wu et al. (2021). It can be assumed that with a further increase in doping, the electronic structure will undergo a Lifshitz transition, significantly reducing the density of states at the Fermi level and, correspondingly, suppressing electronic pairing. This can naturally explain the observed superconducting dome for Tc vs. doping phase diagram Ye et al. (2012); Shi et al. (2015).
@@
Important consequence of this discovery is that the significant part of the total Fermi surface will be spin split at the optimal doping. This agrees with the theoretical prediction Zhao et al. (2020) that with anti-symmetric field effect doping the resulting Fermi surface is dominantly formed by the pockets around $\Sigma$ valleys (only $\sim$30of its area is around ${\rm K}$ valleys as shown in Fig.6(b)). Thus any realistic pairing model of Ising superconductivity in electron doped TMDCs should include such composite spin-split multi-valley Fermi surface topology and corresponding interband scattering processes.
To summarise, we observed spin-polarised Holstein type polaronic quasiparticles at Rb dosed MoSe2 surfaces. This is a direct evidence of the predicted Rashba-like spin-orbit split polarons Covaci and Berciu (2009). We found that beyond the ”surface doping” regime threshold, when dosed alkali atoms intercalate into the van der Waals (vdW) gap between the layers, new plasmon-like quasiparticles emerge. Our findings not only suggest the crucial role of polaron formation for the superconductivity in electron-doped transition-metal dichalcogenides, but also provide experimental foundation for theoretical models of Ising-type superconductivity in these materials.
@@ IV Methods
High quality single crystals of MoSe2 were grown by the chemical vapor transport technique Legma et al. (1993). ARPES measurements were performed at the I05 beamline at the Diamond Light Source, UK Hoesch et al. (2017). The photoelectron’s energy and angular distributions were analysed with a SCIENTA R4000 hemispherical analyser. The angular resolution was 0.2, and the overall energy resolution was better than 5 meV. P-polarised 60 eV photons corresponding to the $\Gamma$ point of the bulk Brillouin zone were used for high-resolution ARPES, and 120 eV photons for Se 3d core level measurements. Single crystal samples were cleaved and measured in ultra high-vacuum at temperatures below 10 K and pressures below 1$\cdot$10-10 mbar. Rb atoms were dosed in several sequences on a cold MoSe2 surface using a commercial SAES alkali-metal dispenser. The density functional theory (DFT) calculations including spin-orbit coupling with generalized gradient approximation (GGA) were performed using WEIN2K package Blaha et al. (2020). Spectral function simulation for polarons has been done by using momentum average (MA) approximation Goodvin et al. (2006) with 20 meV acoustic phonons.
@@ V Acknowledgements
A single 3d5/2/3d3/2 doublet for a pristine MoSe2 crystal splits into two components with Rb dosing as shown in Fig. S4(a). Similar to the case of HfTe2El Youbi et al. (2020) one could fit these core level data assuming bulk (P1) and surface (P2) components due to the electron doping of the MoSe2 surface. Obtained peak area ratio and energy differences (or effective chemical shift) are shown in Fig. S4(b,c). Below critical dosing of 120 s one observes increase of the difference between bulk and surface doublets binding energies with dosing time. This corresponds to the gradual accumulation of Rb atoms, increased electron doping of the MoSe2 surface, and corresponding increase of effective electric filed.
@@
To fit ARPES data, we used constant DFT value for the effective mass $m^{*}=0.54m_{e}$ at $\Sigma$ valley and $k_{F}$ values obtained from MDC fits shown in Fig. S9.

### v2

@@ I Introduction
The resulting polaronic bound state can be described by Fröhlich or Holstein Hamiltonians. The Fröhlich model describes the interaction of electrons with phonons, which leads to the formation of large polarons, whose radii are much larger than the lattice constant. The Holstein model, on the other hand, describes the short-range interaction with phonons, which typically leads to the formation of small polarons, whose radii are of the same order of magnitude as the lattice constant Emin (2013).
@@
Figure 1(a) shows the multi-valley scattering between $\Sigma$ and ${\rm K}$ valleys in the Brillouin zone. The electron scattering vectors $q_{1}$ and $q_{2}$ correspond to either $\Gamma$-${\rm K}$ or $\Gamma$-${\rm M}$ phonon wave vectors. The dominant contribution to electron-phonon coupling arises from the highest energy acoustic phonon mode. This behaviour has been demonstrated both experimentally Zeng et al. (2012) and theoretically Kaasbjerg et al. (2012) for MoS2. As in the case of MoS2 Kaasbjerg et al. (2012); Zhao et al. (2018), for MoSe2 the acoustic phonon modes have essentially flat dispersion over an extended region of $k$-space between the ${\rm M}$ and ${\rm K}$ points Bae et al. (2022). Thus, despite their acoustic nature, these phonon modes with characteristic frequency $\Omega_{ph}$ define the energy scale for electron-phonon coupling and corresponding polaron formation Zhao et al. (2018); Bae et al. (2022). However, only ${\rm K}$ valley Holstein polarons (Fig. 1(b)) have been observed in the ARPES study of Rb-dosed MoS2 surface  Kang et al. (2018). Second, the spin-orbit coupling (SOC) should induce the splitting of the electron bands at $\Sigma$ and $\Sigma$’ valleys due to inversion symmetry breaking on the surface with alkali metal dosing. However, in all previous ARPES studies of alkali-dosed semiconducting TMDCs the spin-split features at $\Sigma$ valley have not been resolved Zhang et al. (2014); Eknapakul et al. (2014); Alidoust et al. (2014); Riley et al. (2015); Miwa et al. (2015); Zhang et al. (2016); Kim et al. (2016, 2017); Kang et al. (2017, 2018); Han et al. (2021).
Figure 2: Evolution of the MoSe2 electronic structure with Rb dosing: (a-a5) ARPES spectra of $\Gamma$-${\rm K}$ conduction band dispersions measured near the Fermi energy for cumulative dosing time 0, 20, 60, 100, 220, 580s. (b-b5) Corresponding valence band spectra in a wide energy range.
The isostructural MoSe2 compound is the perfect model system for probing the multi-valley physics induced by electron-doping. Density functional theory (DFT) calculations show that in contrast to MoS2, the conduction band minimum (CBM) in bulk MoSe2 is at the $\Sigma$, rather than at ${\rm K}$ valley (see Supplementary Fig. S1). For the alkali-metal dosing on MoSe2 surface one expects to have larger SOC band-splitting than in the case of MoS2. The accessibility of the $\Sigma$ valley, as well as the ${\rm K}$ valley expected at higher doping Kim et al. (2016, 2017), makes MoSe2 the best candidate to study the physics of multi-valley polaron modes using high-resolution ARPES.
@@ II.1 Evolution of the electronic band structure with dosing
In our ARPES data shown in Fig. 2 the spectral intensity appears at the Fermi level, marking the occupation of the conduction band and corresponding semiconductor to conductor transition with the Rb dosing on the MoSe2 surface. After initial dosing the maximum spectral weight at the Fermi level is observed around $\Sigma$ valley, showing an indirect band gap $E_{g}^{\Gamma-\Sigma}$ $\sim$1.4 eV. However, with increasing amount of dosed Rb the relative intensity of the $\Sigma$ valley decreases, whereas the intensity of ${\rm K}$ valley increases. Correspondingly, it appears that the conduction band minimum shifts from $\Sigma$ to ${\rm K}$ valley as in the case of monolayer MoSe2 (see Fig. 2(a1-a3)). Withing observed energy broadening the band minima at $\Sigma$ and ${\rm K}$ are very close (see Supplementary Fig. S5), incidentally making electron doped MoSe2 a perfect system for studying multi-valley physics. Simultaneously with the shift of the CBM intensity, the intensity of the valence band compared to the conduction band decreases (see Fig. 2 (b1-b3)). This first crossover at around 40 s of Rb dosing marks the transition from the electron doped bulk to the more “monolayer-like” doped surface conduction band structure Eknapakul et al. (2014); Riley et al. (2015); Kang et al. (2017); Kim et al. (2017).
@@ II.2 Holstein polarons in “surface doping” regime
The measured Fermi surface shows two electron-like pockets at $\Sigma$ valley and single electron-like pocket at ${\rm K}$ valley. The second derivative images of $\Sigma$ and ${\rm K}$ valleys in Fig. 3(c,d) both show well-defined discontinuities of the spectral function separated by 20 meV. Moreover, the photoemission spectra integrated withing a selected momentum range between $\Sigma$ and ${\rm K}$ valleys in Fig. 5(a), show discrete non-dispersing energy levels with the same energy separation (dashed vertical black lines in Fig. 5(b)).
@@ II.3 Spin-splitting at the $\Sigma$ valley in “surface doping” regime
While polarons at the spin-degenerate ${\rm K}$-valley has been observed previously for MoS2 Kang et al. (2018), polarons at the spin-polarised $\Sigma$ valleys are discovered for the first time. The spin-orbit splitting of the electron band dispersion at the $\Sigma$ valley is apparent in ARPES data measured along the $\Gamma$-${\rm K}$ direction. This agrees with theoretical predictions of lifting the spin degeneracy under electron doping and much larger spin-orbit splitting at $\Sigma$ compared to ${\rm K}$ due to distinct orbital character of conduction bands at these valleys Brumme et al. (2015); Clark et al. (2019); Boccuni et al. (2024). Figure 4 shows band dispersions at the $\Sigma$ valley together with the Rashba-like model fit constrained by the band positions from the corresponding momentum distribution curves (MDCs) at the Fermi level.
@@
The increase of the Fermi surface size indicates the gradual electron doping of MoSe2 with Rb dosing. From the same data one could see that the magnitude of band-splitting is also increasing with electron doping. This Rashba-like spin-splitting of the bands comes from inversion symmetry breaking due to the electric field created by electron charge transfer from dosed alkali atoms at the surface. Such inversion symmetry breaking on the surface has been seen at the valence band of semiconducting Kang et al. (2017); Kim et al. (2017) as well as at the conduction band of semimetallic Clark et al. (2019) alkali dosed TMDC materials. Figure 4(f) shows how the magnitude of the observed splitting consistently and monotonically becomes larger as the surface charge increases. This observation aligns perfectly with our understanding that the band splitting is Rashba-like and arises from the combination of spin-orbit coupling with the increasing inversion symmetry-breaking electric field at the surface. Moreover, the constancy of intensity for the two components with dosing rules out the case of superposition of the surface and subsurface as a possible origin of the observed splitting.
@@ III Conclusion
By studying electronic structure evolution of MoSe2 with Rb surface dosing we have identified three different regimes: surface doping, inter-layer intercalation and ordered alkali metal layer formation. Figure 6(a) shows the two dimensional (2D) electron carrier density as a function of the alkali metal dosing time, calculated using the Luttinger’s theorem and experimentally determined sizes of the Fermi surface. In the “surface doping” regime we observe continuous electron doping of MoSe2 with corresponding increase of the Fermi surface size and value of the Rashba-like spin-orbit splitting. In this regime, before interlayer intercalation occurs, Rb-dosing on the surface is comparable to doping induced by the electrostatic gating using ionic liquid. Therefore, the electronic structure on MoSe2 under Rb dosing in “surface doping” regime corresponds to the electronic structure of MoSe2 in ionic liquid gating device under positive gate voltage (electron doping).
@@
While our results show that the increasing surface electric field description is no longer applicable when we dose Rb atoms above 120 s, we could extrapolate a gradual increase of the 2D electron carrier density with the alkali metal dosing time within the ”surface doping” regime up to the superconducting transition. In this case, using extrapolation from our ARPES data we conclude that maximum Tc doping level corresponds to the Fermi surface topology where the electron-like pocket at ${\rm K}$ valley touches the outer electron pocket at $\Sigma$ valley (see Fig. 6(b)). In another words, the maximum Tc corresponds to the Fermi level being located exactly at the conduction band saddle point between ${\rm K}$ and $\Sigma$ valleys. This can lead to a dramatic enhancement of pairing strength, due to strong increase of the momentum resolved electron-phonon coupling constant $\lambda$, like in case of Rb intercalated PtTe2  Wu et al. (2021). It can be assumed that with a further increase in doping, the electronic structure will undergo a Lifshitz transition, significantly reducing the density of states at the Fermi level and, correspondingly, suppressing electronic pairing. This can naturally explain the observed superconducting dome for Tc vs. doping phase diagram Ye et al. (2012); Shi et al. (2015).
@@
An important consequence of this discovery is that a significant part of the total Fermi surface will be spin split at the optimal doping. This agrees with the theoretical prediction Zhao et al. (2020) that with anti-symmetric field effect doping the resulting Fermi surface is dominantly formed by the pockets around $\Sigma$ valleys (only $\sim$30of its area is around ${\rm K}$ valleys as shown in Fig.6(b)). Thus any realistic pairing model of Ising superconductivity in electron doped TMDCs should include such composite spin-split multi-valley Fermi surface topology and corresponding interband scattering processes.
To summarise, we observed, for the first time, spin-polarised Holstein type polaronic quasiparticles on Rb dosed MoSe2 surfaces. This is direct evidence of the predicted Rashba-like spin-orbit split polarons Covaci and Berciu (2009). We found that beyond the “surface doping” regime threshold, when dosed alkali atoms intercalate into the van der Waals (vdW) gap between the layers, completely new plasmon-like quasiparticles emerge. Our findings not only suggest the crucial role of polaron formation for the electronic structure of electron-doped transition-metal dichalcogenides, but also provide experimental foundation for theoretical models of Ising-type superconductivity in these materials.
@@ IV Methods
High quality single crystals of MoSe2 were grown by the chemical vapor transport technique Legma et al. (1993). ARPES measurements were performed at the I05 beamline at the Diamond Light Source, UK Hoesch et al. (2017). The photoelectron’s energy and angular distributions were analysed with a SCIENTA R4000 hemispherical analyser. The angular resolution was 0.2, and the overall energy resolution was better than 5 meV. P-polarised 60 eV photons corresponding to the $\Gamma$ point of the bulk Brillouin zone were used for high-resolution ARPES, and 120 eV photons for Se 3d core level measurements. Single crystal samples were cleaved and measured in ultra high-vacuum at temperatures below 10 K and pressures below 1$\cdot$10-10 mbar. Rb atoms were dosed in several sequences on a cold MoSe2 surface using a commercial SAES alkali-metal dispenser. The density functional theory (DFT) calculations including spin-orbit coupling with generalized gradient approximation (GGA) were performed using Wien2K package Blaha et al. (2020). Spectral function simulation for polarons has been done by using momentum average (MA) approximation Goodvin et al. (2006) with 20 meV constant phonon energy.
@@ V Acknowledgements
A single 3d5/2/3d3/2 doublet for a pristine MoSe2 crystal splits into two components with Rb dosing as shown in Fig. S4(a). Similar to the case of HfTe2 El Youbi et al. (2020) one could fit these core level data assuming bulk (P1) and surface (P2) components due to the electron doping of the MoSe2 surface. Obtained peak area ratio and energy differences (or effective chemical shift) are shown in Fig. S4(b,c). Below critical dosing of 120 s one observes increase of the difference between bulk and surface doublets binding energies with dosing time. This corresponds to the gradual accumulation of Rb atoms, increased electron doping of the MoSe2 surface, and corresponding increase of effective electric field. Beyond this critical dosing time, Rb atoms penetrate the van der Waals gap underneath the top layer, thereby physically and electronically isolating the topmost MoSe2 layer. This clearly marks the second crossover from surface doping to the interlayer intercalation regime.
@@
To fit the ARPES data, we used constant DFT value for the effective mass $m^{*}=0.54m_{e}$ at $\Sigma$ valley and $k_{F}$ values obtained from MDC fits shown in Fig. S10. From these $k_{F}$ values we estimate the experimental asymmetry of the Rashba splitting at $\Sigma$ valley of about $\sim$9%.